# On minimizing cyclists’ ascent times

### v5

@@ 1 Introduction
We begin this article by invoking a phenomenological model to estimate the power required to maintain a given speed. This model is analogous to the one used and discussed by Bos et al. (2023, 2024). Subsequently, we seek the riding strategy that results in the least ascent time, which — under stated constraints and assumptions — we show to be constant speed. We proceed to exemplify this result for several distinct ascents. We conclude by discussing the obtained result, comparing it to another plausible strategy: constant power, and commenting on possible future developments. In the appendices, we provide a method for imposing an optimization constraint, namely, the value of the available average power, justify an aspect of our model simplification, namely, neglecting the power used to increase kinetic energy, examine the effect of imposing the second constraint, namely, the value of the available maximum power, comment on numerical optimization, discuss an ad hoc constant-ascent-speed strategy, and conclude with an insight into empirical adequacy of the model.
@@ 2 Phenomenological model
We consider a phenomenological model for a bicycle-cyclist system of mass $m$ moving with ground speed $V$ [^5] To be consistent with our previous study, (Bos et al., 2023, 2024), where we distinguish between the wheel speed, $v$ , and the centre-of-mass speed, $V$ , herein we use implicitly the latter. In this model, formally, $v\equiv V$ ; however, conceptually, we consider the motion of the centre of mass. The effort expended by the cyclist is quantified by power $P$ , which is composed of power $P_{K}$ , required to change the kinetic energy, $\tfrac{1}{2}mV^{2}$, and the power to overcome opposing forces, $P_{F}$ , as well as to overcome the drivetrain resistance,
@@ 3.2 Critical point
for any pair of segments $i$ and $j$. It follows immediately that $V_{i}=V_{j}$, for any pair of segments. In other words, for a given average power, the critical ascent time is achieved with a constant ground speed regardless of the slope angle. This conclusion is a consequence of the following:
@@ 4.3 Time
[…]
@@ Acknowledgements
We wish to acknowledge Matteo Bertrand, for providing the measurements of his ascents, David Dalton, for his proofreading, Elena Patarini, for her graphic support, and Roberto Lauciello, for his artistic contribution.
@@ Appendix D Numerical optimization considerations
P1. [Initialize.] Set segment speeds to the constant speed, $V_{j}=V$, that corresponds to the global minimum of the ascent time, where the only constraint is the average power, using coefficients (E.1) in solution (18), set $\mathcal{J}=\emptyset$, and initialize a counter $i=1$.
@@
P3. [Update constant speed.] Recalculate $V_{j}=V$ for $j\not\in\mathcal{J}$, using $p$ from expression (C.5) and $q$ from expression (E.1), iterate the counter $i=i+1$, and return to P2.
@@ Appendix F On empirical adequacy
To examine the model, we consider the following values in expression (1) : $m=78.6\,\rm kg$, as reported by the rider, ${\rm C_{d}A}=0.35\,{\rm m}^{2}$, ${\rm C_{rr}}=0.005$, $\lambda=0.02$, as estimated from standard measurements, $\rho=1.1464\,\rm kg/m^{3}$, as calculated based on altitude and temperature, and $g=9.81\,\rm m/s^{2}$. According to the constant-power strategy — with $P_{0}=322\,\rm W$ — $T=541.7489\,\rm s$, which is $9$ minutes and $2$ seconds; the ascent times agree to within one second.

### v6

@@ 1 Introduction
We begin this article by invoking a phenomenological model to estimate the power required to maintain a given speed. This model is analogous to the one used and discussed by Bos et al. (2023, 2024a). Subsequently, we seek the riding strategy that results in the least ascent time, which — under stated constraints and assumptions — we show to be constant speed. We proceed to exemplify this result for several distinct ascents. We conclude by discussing the obtained result, comparing it to another plausible strategy: constant power, and commenting on possible future developments. In the appendices, we provide a method for imposing an optimization constraint, namely, the value of the available average power, justify an aspect of our model simplification, namely, neglecting the power used to increase kinetic energy, examine the effect of imposing the second constraint, namely, the value of the available maximum power, comment on numerical optimization, discuss an ad hoc constant-ascent-speed strategy, give an insight into empirical adequacy of the model, present an analytical Euler-Lagrange solution for optimal ground speed, and conclude with a formulation of an inverse problem to estimate the model parameters by minimizing the discrepancy between model predictions and measurements.
@@ 2 Phenomenological model
We consider a phenomenological model for a bicycle-cyclist system of mass $m$ moving with ground speed $V$ [^5] To be consistent with our previous study, (Bos et al., 2023, 2024a), where we distinguish between the wheel speed, $v$ , and the centre-of-mass speed, $V$ , herein we use implicitly the latter. In this model, formally, $v\equiv V$ ; however, conceptually, we consider the motion of the centre of mass. The effort expended by the cyclist is quantified by power $P$ , which is composed of power $P_{K}$ , required to change the kinetic energy, $\tfrac{1}{2}mV^{2}$, and the power to overcome opposing forces, $P_{F}$ , as well as to overcome the drivetrain resistance,
@@ 3.2 Critical point
for any pair of segments $i$ and $j$. It follows immediately that $V_{i}=V_{j}$, for any pair of segments. In other words, for a given average power, the critical ascent time is achieved with a constant ground speed regardless of the slope angle. This conclusion is a consequence of the following points.
@@ 4.3 Time
In both cases, we calculate the ascent time, $T$, using expression (2) [^8] For the constant-speed case, $T$ can be also obtained analytically, as shown in Appendix G. and tabulate the results in Table 1. $T$ is the shortest for ascent (4(a)) because the distance travelled is the shortest. $T$ is the same for ascents (4(b))–(4(e)) because their distances are equal, as stated in Section 4.1. In Figure 5, we present the power and speed for the constant-speed and constant-power cases for ascents (4(a))–(4(e)).
@@ Acknowledgements
We wish to acknowledge Matteo Bertrand and Alberto Demicheli, guided by G. Andrea Oliveri, for providing the measurements of the ascents, David Dalton, for his proofreading, Elena Patarini, for her graphic support, and Roberto Lauciello, for his artistic contribution. We also wish to acknowledge insightful comments of an anonymous reviewer of Bos et al. (2024b) that led to the derivation presented in Appendix G.
@@ Appendix D Numerical optimization considerations
P1. [Initialize.] Set segment speeds to the constant speed, $V_{j}=V$, that corresponds to the global minimum of the ascent time, where the only constraint is the average power, using coefficients (17) in solution (18), set $\mathcal{J}=\emptyset$, and initialize a counter $i=1$.
@@
P3. [Update constant speed.] Recalculate $V_{j}=V$ for $j\not\in\mathcal{J}$, using $p$ from expression (C.5) and $q$ from expression (17), iterate the counter $i=i+1$, and return to P2.
@@ Appendix F On empirical adequacy
To examine the model, we consider the following values in expression (1) : $m=78.6\,\rm kg$, as reported by the rider, ${\rm C_{d}A}=0.35\,{\rm m}^{2}$, ${\rm C_{rr}}=0.005$, $\lambda=0.02$, which are common values in this context, [^9] Specific estimates of these three model parameters, for another cyclist, are presented in Appendix H. $\rho=1.1464\,\rm kg/m^{3}$, as calculated based on altitude and temperature, and $g=9.81\,\rm m/s^{2}$. According to the constant-power strategy — with $P_{0}=322\,\rm W$ — $T=541.7489\,\rm s$, which is $9$ minutes and $2$ seconds; the ascent times agree to within one second. This specific agreement is obtained by adjusting — within the range of values expected for the rider in question — the three model parameters, ${\rm C_{d}A}$, ${\rm C_{rr}}$ and $\lambda$. This limited range of adjustments ensures that the measurements can refute the conjectured model, in the spirit of Popper’s conjectures and refutations. In other words, an agreement between predictions and measurements supports the empirical adequacy of the model.
@@
Appendix G Calculus of variations approach
The discrete model presented in the main text has its continuous counterpart, wherein the minimum ascent time is achieved in the context of the calculus of variations.
Setting $V(s)=\left.{\rm d}s/{\rm d}t\,\right|_{t=t(s)}$, the Lagrangian function stated in expression (6) is
$\mathcal{L}\left(V,\Lambda\right)=\int\limits_{0}^{T}\frac{{\rm d}s}{V}+\Lambda\left(\frac{1}{T}\int\limits_{0}^{T}\frac{P}{V}\,{\rm d}s-P_{0}\right).$
Multiplying both sides by $T$, we get
$T\,\mathcal{L}=T^{2}+\Lambda\left(\int\limits_{0}^{T}\frac{P}{V}\,{\rm d}s-T\,P_{0}\right).$
Differentiating both sides, we obtain
$\left(\nabla T\right)\mathcal{L}+T\left(\nabla\mathcal{L}\right)=2T\nabla T+\Lambda\left(-\int\limits_{0}^{L}\frac{P}{V^{2}}\nabla V\,{\rm d}s+\int\limits_{0}^{L}\frac{{\rm d}P}{{\rm d}V}\frac{\nabla V}{V}\,{\rm d}s-P_{0}\nabla T\right).$
To obtain a minimum, we set $\nabla\mathcal{L}=0$, for which $\left.\mathcal{L}\right|_{\nabla\mathcal{L}=0}=T$; hence,
$0=\int\limits_{0}^{L}\left((T-\Lambda P_{0})\left(-\frac{1}{V^{2}}\right)-\Lambda\left(\frac{P}{V^{2}}-\frac{1}{V}\frac{{\rm d}P}{{\rm d}V}\right)\right)\nabla V\,{\rm d}s,$
which is a form of the Euler-Lagrange equation. To satisfy this equation for arbitrary integration limits, we require
$0=(T-\Lambda P_{0})\left(-\frac{1}{V^{2}}\right)-\Lambda\left(\frac{P}{V^{2}}-\frac{1}{V}\frac{{\rm d}P}{{\rm d}V}\right),$
which we rearrange to write
$0=\Lambda P_{0}-T-\Lambda\left(P-V\frac{{\rm d}P}{{\rm d}V}\right)\,.$ (G.1)
To state the expression in parentheses within the context of our model, we invoke equation (1), namely,
$P=\dfrac{1}{1-\lambda}\left(m\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{2}\right)V.$
Also,
$\frac{\rm d}{{\rm d}V}\left(V\frac{{\rm d}V}{{\rm d}t}\right)=\frac{{\rm d}V}{{\rm d}V}\frac{{\rm d}V}{{\rm d}t}+V\frac{{\rm d}}{{\rm d}V}\left(\frac{{\rm d}V}{{\rm d}t}\right)=(1)\frac{{\rm d}V}{{\rm d}t}+V\left(\frac{{\rm d}}{{\rm d}t}\left(\frac{{\rm d}V}{{\rm d}V}\right)\right)=\frac{{\rm d}V}{{\rm d}t}+V\left(\frac{{\rm d}}{{\rm d}t}\left(1\right)\right)=\frac{{\rm d}V}{{\rm d}t}\,.$
Hence, the derivative in parentheses in expression (G.1) is
$\frac{{\rm d}P}{{\rm d}V}$ $=\frac{\rm d}{{\rm d}V}\left(\dfrac{1}{1-\lambda}\left(mV\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)V+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}\right)\right)$
$=\frac{1}{1-\lambda}\left(\dfrac{\rm d}{{\rm d}V}\left(mV\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)V+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}\right)\right)$
$=\frac{1}{1-\lambda}\left(m\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)+\frac{3}{2}{\rm C_{d}A}\,\rho\,V^{2}\right)\,,$
which results in the second term therein,
$V\frac{{\rm d}P}{{\rm d}V}=\frac{1}{1-\lambda}\left(\left(m\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)\right)V+\frac{3}{2}{\rm C_{d}A}\,\rho\,V^{3}\right)\,.$
Subsequently,
$\left(1-\lambda\right)\left(P-V\frac{{\rm d}P}{{\rm d}V}\right)$ $=\left(\left(m\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)\right)V+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}\right)$
$\qquad-\left(\left(m\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)\right)V+\frac{3}{2}{\rm C_{d}A}\,\rho\,V^{3}\right)$
$=-{\rm C_{d}A}\,\rho\,V^{3}\,.$
Thus, the Euler-Lagrange equation, stated in expression (G.1), becomes
$0=\Lambda P_{0}-T-\Lambda\left(-{\rm C_{d}A}\,\rho\,V^{3}\right)\,,$
which implies that
${\rm C_{d}A}\,\rho\,V^{3}=\frac{T}{\Lambda}-P_{0}\,,$
and therefore $V$ is a constant.
To find that constant, we must satisfy the average power constraint,
$\frac{1}{T}\int\limits_{0}^{T}P\,{\rm d}t=P_{0}\,.$
Since ${\rm d}s/{\rm d}t=V\implies{\rm d}t={\rm d}s/V$,
$\frac{1}{T}\int\limits_{0}^{L}P\,\frac{{\rm d}s}{V}=P_{0}\,.$
Since for a constant speed, $T=L/V$,
$\frac{V}{L}\int\limits_{0}^{L}P\,\frac{{\rm d}s}{V}=\frac{1}{L}\int\limits_{0}^{L}P\,{\rm d}s=P_{0}\,.$
Returning to equation (1),
$P=\dfrac{1}{1-\lambda}\left(m\dfrac{{\rm d}V}{{\rm d}t}+mg\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{2}\right)V,$
with ${\rm d}V/{\rm d}t=0$, for a constant $V$, we write
$\left(1-\lambda\right)\int\limits_{0}^{L}P\,{\rm d}s$ $=\int\limits_{0}^{L}\left(m\left(0+g\left(\sin\theta+{\rm C_{rr}}\cos\theta\right)\right)V+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}\right){\rm d}s$
$=m\,g\,V\int\limits_{0}^{L}\left(\sin\theta+{\rm C_{rr}}\cos\theta\right){\rm d}s+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}\int\limits_{0}^{L}\left(1\right){\rm d}s$
$=m\,g\,V\int\limits_{0}^{L}\left(\sin\theta+{\rm C_{rr}}\cos\theta\right){\rm d}s+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}L$
If the ascent profile is given by $\left(x(s),y(s)\right)$, where $0\leqslant s\leqslant L$, and parameterized by arclength, $\left(x^{\prime}(s),y^{\prime}(s)\right)=\left(\cos\theta,\sin\theta\right)$, the unit tangent vector is
$\cos\theta=x^{\prime}(s)\quad\text{and}\quad\sin\theta=y^{\prime}(s).$
Hence
$\int\limits_{0}^{L}\left(\sin\theta+{\rm C_{rr}}\cos\theta\right){\rm d}s=\int\limits_{0}^{L}\left(y^{\prime}(s)+{\rm C_{rr}}x^{\prime}(s)\right){\rm d}s=\left(y(L)-y(0)\right)+{\rm C_{rr}}\left(x(L)-x(0)\right)=H+{\rm C_{rr}}\,D,$
where $H:=y(L)-y(0)$ is the height gain and $D:=x(L)-x(0)$ is the horizontal gain.
It follows that
$\left(1-\lambda\right)\int\limits_{0}^{L}P\,{\rm d}s=m\,g\,V\left(H+{\rm C_{rr}}\,D\right)+\frac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}L$
and thus the average power constraint is
$\frac{1}{L}\int\limits_{0}^{L}P\,{\rm d}s=P_{0}\iff\frac{m\,g\,V\left(H+{\rm C_{rr}}\,D\right)}{(1-\lambda)L}+\dfrac{\tfrac{1}{2}{\rm C_{d}A}\,\rho\,V^{3}}{1-\lambda}=P_{0}\,.$
This is a cubic equation for $V$, which can be solved numerically; there is only one positive real solution. There is also the somewhat laborious, yet analytical, Cardano’s solution.
To exemplify this formulation, let us consider $m=70\,{\rm kg}$, ${\rm C_{d}A}=0.3\,{\rm m^{2}}$, ${\rm C_{rr}}=0.005$, $\lambda=0.02$, $g=9.81\,{\rm m/s^{2}}$, $H=100\,{\rm m}$, $D=1000\,{\rm m}$, $L=1004.9876\,{\rm km}$, $\rho=1.2\,{\rm kg/m^{3}}$ and $P_{0}=300\,{\rm W}$, which results in
$V^{3}+398.5870V=1633.3333\,.$
Its only real solution is $V=3.9439\,{\rm m/s}$, as expected in view of the first row in Table 1. Letting $L=1006.6272\,{\rm m}$, we obtain a slightly different cubic whose only real solution is the speed for the constant-speed strategy in the remaining rows of that table. In all cases, the corresponding ascent times are $T=L/V$
The constant-power case leads to a first order nonlinear differential equation for $V(t)$ ; hence, in contrast to the constant-speed case presented herein, it has no explicit solution.
Appendix H On estimation of resistance parameters
In Appendix F, we demonstrate the empirical adequacy of power model (1) through its agreements with measurements of ascent time and instantaneous speeds of a cyclist along the San Bernardo ascent. These agreements depend on accurate estimates of the air, ${\rm C_{d}A}$, rolling, ${\rm C_{rr}}$, and drivetrain, $\lambda$, resistance parameters, which we obtain by adjusting their values within the expected range. In this appendix, we demonstrate a procedure to estimate these values using numerical optimization.
For this purpose, we obtain further measurements along the San Bernardo ascent; these ascents are done by another cyclist. Following the approach of Appendix F, the cyclist used a flying start and the constant-power strategy. For a statistical evaluation, the cyclist repeated the ascent three times, striving to maintain the same power along a given ascent.
The data are collected using a Garmin 1030 cyclocomputer in combination with pedal-based 4iii power metres. The data are sampled once a second and measure altitude [m], ambient temperature [${}^{\circ}{\rm C}$], cadence [rpm], distance [m], heartrate [bpm], power [W], time [s], and velocity [m/s]. The environmental conditions during the efforts correspond to a windless and sunny day with a temperature of approximately $22^{\circ}{\rm C}$. Using the temperature, along with the altitude, we estimate the air density, $\rho=1.1565\,{\rm kg}/{\rm m}^{3}$, using the formula presented in Danek et al. (2021, Section 5.3). Measurements of ascent time, $\mathcal{T}$, average power, $\overline{\mathcal{P}}$, and distance travelled, $L$, are summarized in Table 2 for each effort. We note that — in spite of the fact that each ascent corresponds to the same Strava [^10] Strava is an internet service for tracking physical exercise using GPS. segment — the distance travelled is different for each effort, which is due to the rather sparse sampling of the cyclocomputer and the position error of the GPS.
Data measurements Optimized resistance parameters
effort $\mathcal{T}\,{\rm[s]}$ $\overline{\mathcal{P}}\,{\rm[W]}$ $L\,{\rm[m]}$ ${\rm C_{d}A}\,[{\rm m}^{2}]$ ${\rm C_{rr}}$ $\lambda$
1 $579$ $247\pm 13$ $3486$ $0.3521\pm 0.0128$ $0.0059\pm 0.0005$ $0.0317\pm 0.0017$
2 $577$ $247\pm 15$ $3426$ $0.3671\pm 0.0102$ $0.0062\pm 0.0004$ $0.0321\pm 0.0016$
3 $489$ $312\pm 27$ $3467$ $0.3575\pm 0.0084$ $0.0059\pm 0.0004$ $0.0313\pm 0.0010$
Table 2: Optimization of resistance parameters for three efforts of the San Bernardo ascent following the constant-power strategy
To estimate the resistance parameters in model (1), we require that ${\rm C_{d}A}$, ${\rm C_{rr}}$, and $\lambda$ result in a modelled ascent time, $T$, that agrees with the measured ascent time, $\mathcal{T}$. In other words, we seek to obtain numerical solutions to $f({\boldsymbol{\beta}})=\left(T({\boldsymbol{\beta}})-\mathcal{T}\right)^{2}$, where ${\boldsymbol{\beta}}=\left[\,{\rm C_{d}A},\,{\rm C_{rr}},\,\lambda\,\right]^{T}$, such that the power along each segment of an effort is fixed to the measured average power, $\overline{\mathcal{P}}$. We solve the problem using MATLAB (2022)’s fmincon solver, which finds the minimum of a constrained nonlinear multivariable function. To maintain reasonable estimates, we apply bounds to the resistance parameters in the optimization such that ${\rm C_{d}A}\in[0.1,0.5]\,{\rm m}^{2}$, ${\rm C_{rr}}\in[0.001,0.01]$, $\lambda\in[0.01,0.05]$. Also, in contrast to the cyclist in Appendix F, the mass of the bicycle-cyclist system is $m=60.6\,{\rm kg}$.
Given the underdetermined nature of the problem, there are many combinations of ${\rm C_{d}A}$, ${\rm C_{rr}}$ and $\lambda$ that result in an agreement between modelled and measured ascent time, $T=\mathcal{T}$, under the assumption of constant power, $P_{j}=\overline{\mathcal{P}}$ , for $j=1,\dots,N$. For this reason, we repeat the optimization 10 0001000010\,00010 000 times for each effort using different starting resistance values sampled uniformly from the specified parameter bounds. The sample mean and standard deviation of the optimized resistance parameters are stated in Table 2. We see that the resistance-parameter estimates are similar for each of the three ascents, which is a supporting evidence of empirical adequacy of the model.